OpenStudy (therealmeeeee):

help given medals Which value for x makes the open sentence true? 5 + 4 • x = 32 + x2 A. 2 B. 3 C. 4 D. 5

OpenStudy (loser66):

x2 means?

OpenStudy (therealmeeeee):

times 2

OpenStudy (loser66):

and what does 4 . x mean?

OpenStudy (therealmeeeee):

idk

OpenStudy (loser66):

hahaha.... decimal number like 0. 5 ?

OpenStudy (therealmeeeee):

so it 5

OpenStudy (loser66):

nope, I just want to confirm whether I understand the problem or not.

OpenStudy (therealmeeeee):

i have 5 question

OpenStudy (therealmeeeee):

do u

OpenStudy (therealmeeeee):

@JWOWW_98

OpenStudy (loser66):

solve this one first, please. You have to know what they mean to solve . 4* x \(\neq\) 4.x, right? 2x \(\neq x^2\), right?

OpenStudy (therealmeeeee):

wat is that

OpenStudy (loser66):

@jim_thompson5910

OpenStudy (therealmeeeee):

@JonnyVonny

OpenStudy (therealmeeeee):

@dan815

OpenStudy (loser66):

@jdoe0001

OpenStudy (therealmeeeee):

@TheSecretsToday

OpenStudy (therealmeeeee):

@SeaTurtle113

OpenStudy (therealmeeeee):

@Luigi0210

OpenStudy (loser66):

@amistre64

OpenStudy (therealmeeeee):

plzzzz help me

OpenStudy (therealmeeeee):

i have 5 questions

OpenStudy (anonymous):

this is actually pretty easy

OpenStudy (anonymous):

have u tried plugging in the numbers

OpenStudy (therealmeeeee):

yes

OpenStudy (therealmeeeee):

can u just help me

OpenStudy (therealmeeeee):

omg i need helpppppppppppppppppppppp

OpenStudy (therealmeeeee):

with 5 questions

OpenStudy (therealmeeeee):

@lulu22

OpenStudy (jonnyvonny):

@TheSecretsToday , from a first glance at it, if x is to the power of 2, it is impossible; @TheRealMeeeee , may you please rewrite the equation as a post, so it may be more accurate?

OpenStudy (therealmeeeee):

watttttttttttttttttttt

OpenStudy (therealmeeeee):

some one plzzzz help me with 5 question

OpenStudy (luigi0210):

Have you consider posting one question at a time? ლ(ಠ_ಠლ)

OpenStudy (therealmeeeee):

ots is one

OpenStudy (therealmeeeee):

im on question 1

OpenStudy (therealmeeeee):

ლ(ಠ_ಠლ)

OpenStudy (luigi0210):

Oh, you're doing PEMDAS huh?

OpenStudy (therealmeeeee):

i thinl so

OpenStudy (therealmeeeee):

so dont none of u no how to do this

OpenStudy (luigi0210):

I know @Loser66 knows how to do this.. just clarify things. I gtg take care of some stuff. Good luck. ლ(ಠ_ಠლ)

OpenStudy (therealmeeeee):

ok

OpenStudy (therealmeeeee):

@Loser66 may u help me plzzzz

OpenStudy (therealmeeeee):

ლ(ಠ_ಠლ)

OpenStudy (jonnyvonny):

@TheRealMeeeee Type our the equation again so we may see it.

OpenStudy (therealmeeeee):

Which value for x makes the open sentence true? 5 + 4 • x = 32 + x2 A. 2 B. 3 C. 4 D. 5

OpenStudy (therealmeeeee):

ლ(ಠ_ಠლ)

OpenStudy (amistre64):

5 + 4x = 32 + x^2 its a quadratic equation :/

OpenStudy (therealmeeeee):

yeah i think

OpenStudy (therealmeeeee):

@amistre64

OpenStudy (amistre64):

then use the qudratic formula to find the solutions

OpenStudy (therealmeeeee):

nvm ill do that one later on